# Survey on Large Language Model-Enhanced Reinforcement Learning: Concept, Taxonomy, and Methods

### v1

@@ IV-A2 Fine-tuning Pre-trained Model
Some of the other literature modifies the pre-trained models with contrastive learning to learn an invariant feature representation, which improves the generalization capability.
Contrastive learning is a method that aims to learn representations from high-dimensional data by contrasting positive examples against negatives. Given a query $q$, the goal is to match query $q$ more closely to a positive key $k_{+}$ than to any negative keys $\mathbb{K}\setminus\{k_{+}\}$ in a set $\mathbb{K}$. This process is modeled using similarity measures, such as the dot product $(q^{T}k)$ or the bilinear product $(q^{T}Wk)$, where $W$ is a weight matrix. To effectively learn these representations, a contrastive loss function such as InfoNCE [78] is used:
$\mathcal{L}_{q}=\log\frac{\exp(q^{T}Wk_{+})}{\exp(q^{T}Wk_{+})+\sum_{k}^{K-1}\exp(q^{T}Wk_{i})}$ (1)
where $\exp$ is the exponential symbol and the whole $\mathcal{L}_{q}$ can be viewed as the log-loss of a softmax classifier, treating the matching of $q$ to $k_{+}$ as a multi-class classification problem among $K$ classes. By maximizing alignment between different changes of the same observation via the above loss, the model can learn the invariant representations, which addresses the adaptation problem induced by the variations of real-world environments and improves the sample efficiency in out-of-distribution generalization.
In reference [79], to achieve the zero-shot capability of embodied agents, the author devises a visual prompt-based contrastive learning framework that uses a pre-trained VLM to learn the visual state representations. By contrastively training the VLM on a pool of visual prompts along with the RL policy learning process, the learned representations are robust to the variations of environments and achieve competitive zero-shot capabilities. Based on the contrastive learning, another method ReCoRe [80] adds an intervention-invariant regularizer in the form of an auxiliary task such as depth prediction and image denoising to explicitly enforce invariance of learned representations to environment changes.
@@ IV-B1 Instruction Information Translation
One application of LLM is to translate natural language-based instructions for instruction-following applications. In reference [43], the author investigates an inside-out scheme for natural language-conditioned RL by training an LLM that translates the natural language to a task-related unique language. Such an inside-out scheme prevents the policy from being directly exposed to natural language instructions and helps efficient policy learning. Another literature STARLING [81] uses LLM to translate natural language-based instructions to game information and example game metadata. The translated information is then fed forward to Inform7, an interactive fiction game engine, to develop a large amount of text-based games for RL agents to master the desired skills.
@@ IV-C Summarization and Outlook
For information processing, LLM is used to accelerate RL learning processing by decoupling the information processing task and the controlling task, where LLM extracts feature representations or handles the natural language-based information. In the following, we list potential directions for future research.
  • Feature Representation Extractor: Although the use of LLMs as feature representation extractors has shown promise in enhancing RL, several challenges persist in future research. First, the computational overhead of LLMs is large for RL agents, an efficient and simplified feature extractor without excessive computational demand is required. Second, although LLMs can be used to learn features from high-dimensional inputs, their representation may lack generalization and be vulnerable to the high diversity inherent in environments or not take into account specific characteristics for solving control tasks. Future work may consider exploiting the task compositionality to increase the generalization of learned representations.
  • Language Translator: Current researches only focus on language translation and directly trains LLM to translate the natural language to task-related language while not considering the more generalization problem where the environment or task information is multi-modal. Future work may consider adding the multi-modal translation ability to achieve a more general translator and help agent learning.
@@ V LLM as Reward Designer
The reward signal is the most important information to instruct agent learning in RL [22]. However, despite the fundamental importance, high-performing reward functions are known to be notoriously difficult to design [87]. On the one hand, specifying human notions of desired behavior is difficult via designed reward functions or requires huge expert demonstrations. On the other hand, dense rewards that accurately provide learning signals require either manually decomposing the general goal into sub-goals [88] or rewarding interesting auxiliary objectives [89]. Nevertheless, both of these methods suffer from the need for expert input and meticulous manual crafting [23].
@@
Figure 4: LLM as a reward designer.
@@ V-A1 Direct Prompting
Reference [90] simplifies the manual reward design by prompting an LLM as a proxy reward function with examples of desirable behaviors and the preferences description of the desired behaviors. Reference [91] proposes a Read and Reward framework that utilizes LLMs to read instruction manuals to boost the learning policies of specific tasks. The framework includes a QA extraction module for information retrieval and summarization and a Reasoning module for evaluation. Experimentally, they show RL algorithms by their design can obtain significant improvement in performance and training speed. In reference [39], Carta et al. proposes an automated reward shaping method where the agent extracts auxiliary objectives from the general language goal. Using a question generation (QG) and question answering system (QA), the framework guides the agent in reconstructing partial information about the global goal and provides an intrinsic reward signal for the agent. This intrinsic reward incentivizes the agent to produce trajectories that help them reconstruct the partial information about the general language goal. To acquire a generalizable policy by continually learning a set of tasks is challenging for RL agents since the agent is required to retain the previous knowledge while quickly adapting to new tasks. Reference [92] introduces the Lafite-RL (Language agent feedback interactive Reinforcement Learning) framework that provides interactive rewards mimicking human feedback based on LLMs’ real-time understanding of the agent’s behavior. By designing two prompts, one to let LLM understand the scenario and the other to instruct it about the evaluation criterion, their framework can accelerate the RL process while freeing up human effort during the interaction between agent and environment.
@@ V-A2 Alignment Scoring
For visual RL, some literature utilizes vision-language models as reward models to align multi-modal data and calculate the similarity score using metrics such as cosine similarity. Rocamonde et al. employs the CLIP model as a zero-shot reward model to specify tasks via natural language [53]. They first compute the probability $p_{o_{t},l}$ that the agent achieves a goal given by language description $l$ out of a set of potential goals $l^{\prime}\in\mathcal{L}$ in the task set $\mathcal{L}$ using the softmax computation with temperature $\tau$ over the cosine similarity between visual state embeddings $f_{\theta}(o_{t})$ and language description embeddings $g_{\theta}(l)$ across the set of potential goals $l^{\prime}$:
$p_{o_{t},l}=\frac{\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l)/\tau)}{\sum_{l^{\prime}}\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l^{\prime})/\tau)}$ (2)
Then the reward is obtained by a binary reward function $r_{t}=r(o_{t+1},l)=\mathbb{I}[p_{o_{t+1},l}\geq\beta]$, which thresholding the probability. Their framework only requires a single-sentence text prompt description of the desired task with minimal prompt engineering. Another work [93] constructs reward signals based on the similarity between natural language-based description and embeddings from the pre-trained VLM encoder. By labeling the expert demonstration with the reward signals, the framework effectively mitigates the problem of mis-generalization. In reference [94], the authors propose a Language Reward Modulated Pretraining (LAMP) framework by leveraging the zero-shot capabilities of vision-language models to generate noisy, albeit shaped exploration rewards. Such rewards are generated by computing the contrastive alignment between language instructions and the image observations of the agent in the pre-training environment. Using the LAMP framework, the generated rewards are optimized with standard novelty-seeking exploration rewards for language-conditioned policy and can warm start sample-efficient learning on robot manipulation tasks.
@@ V-B Explicit Reward Model
To help robots learn low-level actions, reference [95] harnesses the code generation ability of LLMs to define the lower-level reward parameters based on high-level instructions. Using such a reward design paradigm, their work bridges the gap between high-level language instructions to low-level robot actions and can reliably tackle 90% of the designed tasks compared to the 50% of the baseline. Motivated by the capability of LLM for self-refinement [96], reference [97] proposes a framework with a self-refinement mechanism for automated reward function design, including initial design, evaluation and self-refinement loop. Their results indicate that the LLM-designed reward functions are able to rival or surpass manually designed reward functions. Similarly, Eureka [98] develops a reward optimization algorithm with self-reflection. The algorithm uses an environment source code and task description to sample reward function candidates from a coding LLM, evaluate them over RL simulation, and use the result to progressively improve the reward outputs. The results show that their proposed method can achieve human-level performance on reward design and solve dexterous manipulation tasks that were previously infeasible by manual reward engineering. Another work Text2Reward [99] generates shaped dense reward functions as executable programs based on the environment description. Given the sensitivity of RL training and the ambiguity of language, the RL policy may fail to achieve the goal. Text2Reward addresses the problem by executing the learned policy in the environment, requesting human feedback and refining the reward accordingly.
@@ V-C Summarization and Outlook
The use of LLMs as reward designers in reinforcement learning offers a more natural and efficient way to design complex reward functions. By leveraging natural language processing capabilities, LLMs simplify the traditionally challenging task of reward function design, enhancing both the efficiency and effectiveness of RL algorithms. In the future, we expect advancements in the following areas:
  • Implicit Reward Model: Future work may consider focusing on improving the alignment of LLM-generated rewards with complex human intentions and behaviors. This involves refining the quality of language instructions to reduce ambiguities and inaccuracies, ensuring that the reward functions align precisely with human notions of desired behavior. Additionally, generalization and transferability of LLM-generated rewards across different tasks and environments, especially in complex, high-dimensional visual environments, is also an important direction to explore.
  • Explicit Reward Model: A key limitation in reward code generation is its dependency on pre-trained common-sense knowledge, which can be restrictive for highly specialized tasks not covered in training data. Enhancing prompts with detailed, task-specific information is essential for effectiveness. Additionally, the reliance on manually designed templates for motion descriptions limits the adaptability, suggesting a need for future research focused on automating or unifying template design processes.
@@ VI LLM as Decision-Maker
LLMs trained on a massive amount of data show impressive results in language understanding tasks [56], instruction-following [100], vision-language navigation [101] and tasks requiring planning and sequential reasoning [102]. Such success motivates researchers to explore the potential of LLMs for decision-making problems as a direct decision-maker that generates actions or an indirect decision-maker that instructs the actions. In direct decision-making, Transformer-based models such as Decision Transformer [103] show great potential in offline RL when dealing with sparse-reward and long-horizon tasks [104]. As LLM itself is a large-scale Transformer-based model, LLM can be viewed as an enhanced Decision Transformer with more pre-trained world knowledge and powerful modeling abilities, which unleash the power of pre-trained LLMs for offline RL [105]. In indirect decision-making, LLMs can be prompted to generate “expert” actions that instruct the action selection of RL agents.
Figure 5: LLM as a decision-maker.
VI-A Direct Decision-Maker
In direct decision-making, LLM enhances the Decision Transformer-based methods with a more powerful pre-trained model with more prior world knowledge to solve sparse-reward and long-horizon tasks with a higher sample efficiency.
As shown in Fig. 5 (i), Decision Transformer-based methods directly predict the actions from the perspective of sequence modeling by framing the learning process of offline RL into a supervised learning problem. Given each trajectory $\tau$ as a sequence of ordered return-to-go $\hat{R}$, action $a$, and states $s$ as follows:
$\tau=($ $\hat{R}_{t_{0}},s_{t_{0}},a_{t_{0}},\hat{R}_{t_{0}+1},s_{t_{0}+1},a_{t_{0}+1},\ldots,$
$\hat{R}_{t_{0}+K-1},s_{t_{0}+K-1},a_{t_{0}+K-1})$ (3)
where $K$ is the context length, the model learns to predict the future action $a^{\prime}_{t}$ by minimizing the squared error loss:
$\mathcal{L}_{decision}=\sum_{t=t_{0}}^{t_{0}+K-1}\left\|a_{t}-a_{t}^{\prime}\right\|_{2}^{2}$ (4)

### v3

@@ IV-A2 Fine-tuning Pre-trained Model
When trained RL agents are deployed in real-world applications, their performance often deteriorates under significant appearance variations (out-of-distribution data) due to overfitting to training scenarios. For example, robots with vision-based navigation tasks may fail when the environment color changes. Invariant feature representations serve as a form of state abstraction that remains consistent across out-of-distribution appearance variations such as added noise, brightness changes, or slight rotations. When encountering appearance changes or out-of-distribution data, though the observation changed, the representation (feature embedding) that fed into the policy/value network is nearly unchanged, leading to robust RL performance and increased generalization in unseen environments.
Contrastive learning is a common way to learn the invariant feature representation. It learns representations from high-dimensional data by contrasting positive examples against negatives. Given a query $q$, the goal is to match query $q$ more closely to a positive key $k_{+}$ than to any negative keys $\mathbb{K}\setminus\{k_{+}\}$ in a set $\mathbb{K}$. This process is modeled using similarity measures, such as the dot product $(q^{T}k)$ or the bilinear product $(q^{T}Wk)$, where $W$ is a weight matrix. To effectively learn these representations, a contrastive loss function such as InfoNCE [91] is used:
$\mathcal{L}^{c}=\log\frac{\exp(q^{T}Wk_{+})}{\exp(q^{T}Wk_{+})+\sum_{k}^{K-1}\exp(q^{T}Wk_{i})}$ (1)
where $\exp$ is the exponential symbol and the whole $\mathcal{L}^{c}$ can be viewed as the log-loss of a softmax classifier, treating the matching of $q$ to $k_{+}$ as a multi-class classification problem among $K$ classes. By maximizing alignment between different changes of the same observation via the above loss, the model can learn the invariant representations.
When combining with RL, given different RL tasks, the required invariant feature representations should be adjusted accordingly. Therefore, researchers have explored different ways to improve contrastive learning. In reference [92], to achieve the zero-shot capability of embodied agents, the author devised a visual prompt-based contrastive learning framework that uses a pre-trained VLM to learn the visual state representations. The visual prompts are learned on expert demonstrations from domain factors such as camera settings and stride length. By contrastively training the VLM on a pool of visual prompts along with the RL policy learning process, the learned representations are robust to the variations of environments, leading an increase of 18-20% success rates on unseen scenarios and improved generalization capability. Based on the contrastive learning, another method ReCoRe [93] added an intervention-invariant regularizer in the form of an auxiliary task such as depth prediction and image denoising to explicitly enforce invariance of learned representations to environment changes.
@@ IV-B1 Instruction Information Translation
One application of LLM is to translate natural language-based instructions for instruction-following applications. In reference [49], the author investigated an inside-out scheme for natural language-conditioned RL by training an LLM that translates the natural language to a task-related unique language. Such an inside-out scheme prevents the policy from being directly exposed to natural language instructions and helps efficient policy learning. Another literature STARLING [94] used LLM to translate natural language-based instructions to game information and example game metadata. The translated information is then fed forward to Inform7, an interactive fiction game engine, to develop a large amount of text-based games for RL agents to master the desired skills.
@@ IV-C Summarization and Outlook
For information processing, LLM is used to accelerate RL learning processing by decoupling the information processing task and the controlling task, where LLM extracts feature representations or handles the natural language-based information.
When multimodal data are involved in the environment, e.g., robot manipulation tasks, the information processing task becomes more challenging. For one thing, the misalignment or contradiction between different modalities may exist [100]. Modality weighting techniques that adaptively learn the importance of different modalities by attention mechanisms provide a potential solution for this problem [101]. In addition, CLIP-based multimodal foundation models using large-scale image-text pairs have shown outstanding zero-shot ability in various multimodal tasks [89]. By learning cross-modal and task semantics, the misaligned modality can be replaced with a virtually generated modality [102]. For another, how to effectively combine the information between different modalities into a unified representation, i.e., multimodal fusion, is also an important issue. In reference [103], a multimodal contrastive learning with attention mechanisms, which learns the intra and inter-modal representations, was proposed. The different attention heads align the agreement from one modality to another and vice versa, providing an informative representation for the RL agent. However, most multimodal learning methods do not consider the task information and only focus on the modality alignment, remaining an area to be explored.
In the following, we list potential directions for future research.
  • Feature Representation Extractor: Although the use of LLMs as feature representation extractors has shown promise in enhancing RL, several challenges persist in future research. Short-term goals include developing a computationally efficient feature extractor and improving the generalization of LLM-derived representations. In the long term, researchers should focus on exploiting task compositionality for better generalization and creating adaptive extraction methods for diverse control tasks.
  • Language Translator: Current existing works are still limited. Short-term objectives involve exploring LLMs’ ability to handle more tasks and improving translation efficiency and accuracy in RL contexts. Long-term goals include developing multimodal translation capabilities and integrating these with RL algorithms, achieving a more general translator, and helping agent learning.
@@ V LLM as Reward Designer
The reward signal is the most important information to instruct agent learning in RL [1]. However, despite the fundamental importance, high-performing reward functions are known to be notoriously difficult to design [104]. First, specifying human notions of desired behavior is difficult via designed reward functions or requires huge expert demonstrations. Moreover, dense rewards that accurately provide learning signals require either manually decomposing the general goal into sub-goals [105] or rewarding interesting auxiliary objectives [106]. Nevertheless, both of these methods suffer from the need for expert input and meticulous manual crafting [23].
@@
Figure 4: LLM as a reward designer. (i) Implicit Reward Model: LLMs provide rewards through direct prompting or alignment scoring between language instructions and visual observations. (ii) Explicit Reward Model: LLMs generate executable code for reward functions, with potential for self-refinement through evaluation loops.
@@ V-A1 Direct Prompting
Reference [107] simplified the manual reward design by prompting an LLM as a proxy reward function with examples of desirable behaviors and the preferences description of the desired behaviors. Reference [108] proposed a Read and Reward framework that utilizes LLMs to read instruction manuals to boost the learning policies of specific tasks. The framework includes a Question & Answer (QA) extraction module for information retrieval and summarization and a Reasoning module for evaluation. Experimentally, they show RL algorithms, by their design, can obtain significant improvement in performance and training speed. In reference [45], Carta et al. proposed an automated reward shaping method where the agent extracts auxiliary objectives from the general language goal. Using a question generation and QA system, the framework guides the agent in reconstructing partial information about the global goal and provides an intrinsic reward signal for the agent. This intrinsic reward incentivizes the agent to produce trajectories that help them reconstruct the partial information about the general language goal. To acquire a generalizable policy by continually learning a set of tasks is challenging for RL agents since the agent is required to retain the previous knowledge while quickly adapting to new tasks. Reference [109] introduced the Lafite-RL (Language agent feedback interactive RL) framework that provides interactive rewards mimicking human feedback based on LLMs’ real-time understanding of the agent’s behavior. By designing two prompts, one to let LLM understand the scenario and the other to instruct it about the evaluation criterion, their framework can accelerate the RL process while freeing up human effort during the interaction between agent and environment.
Algorithm 1 Language Reward Modulated Pretraining (LAMP) in [110]
1:  Initialize parameters for Masked World Models (MWM)
2:  Load pretrained DistilBERT [111] for language preprocessing
3:  Load pretrained R3M visual encoder and score predictor
4:  Initialize empty replay buffer
5:  Populate language prompt buffer and synonym buffers with predefined samples
6:  for each training episode do
7:     Randomize scene textures from Ego4D and RLBench datasets
8:     Sample ShapeNet objects and language prompts
9:     Insert ShapeNet objects into the scene
10:     Generate and process language embeddings using DistilBERT
11:     Execute policy to collect transitions
12:     Assign LAMP rewards using R3M score predictos
13:     Update buffers and train MWM with augmented rewards
14:  end for
@@ V-A2 Alignment Scoring
For visual RL, some literature utilizes vision-language models as reward models to align multimodal data and calculate the similarity score using metrics such as cosine similarity. Rocamonde et al. employs the CLIP model as a zero-shot reward model to specify tasks via natural language [64]. They first compute the probability $p_{o_{t},l}$ that the agent achieves a goal given by language description $l$ out of a set of potential goals $l^{\prime}\in\mathcal{L}$ in the task set $\mathcal{L}$ using the softmax computation with temperature $\tau$ over the cosine similarity between visual state embeddings $f_{\theta}(o_{t})$ and language description embeddings $g_{\theta}(l)$ across the set of potential goals $l^{\prime}$:
$p_{o_{t},l}=\frac{\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l)/\tau)}{\sum_{l^{\prime}}\exp(f_{\theta}(o_{t})\cdot g_{\phi}(l^{\prime})/\tau)}.$ (2)
Then the reward is obtained by a binary reward function $r_{t}=r(o_{t+1},l)=\mathbb{I}[p_{o_{t+1},l}\geq\beta]$, which thresholding the probability. Their framework only requires a single-sentence text prompt description of the desired task with minimal prompt engineering. Another work [112] constructed reward signals based on the similarity between natural language-based description and embeddings from the pre-trained VLM encoder. By labeling the expert demonstration with the reward signals, the framework effectively mitigates the problem of mis-generalization. In reference [110], the authors proposed the Language Reward Modulated Pretraining (LAMP) framework as a pertaining utility for RL as opposed to a downstream task reward to warm-start sample-efficient learning. The framework leverages frozen, pre-trained VLMs such as R3M [113] to generate noisy, albeit shaped exploration rewards by computing the alignment score between instructions and image observations. The algorithm is presented in Algorithm 1. They used Masked World Model [114], a visual model-based RL algorithm for robot manipulation based on image and instructions. The images were downloaded from Ego4D [115] and the language instructions were obtained by querying ChatGPT. The reward is then calculated from the R3M alignment score. After that, the generated rewards are optimized with standard novelty-seeking exploration rewards for language-conditioned policy. Reference [116] explored preference-based RL, where the agent learns a reward function from preference labels over the behaviors. A VLM is leveraged to generate preference labels given visual observations and a text description of the task goal. The evaluation is based on a series of vision-based manipulation tasks. Results suggested that prompting VLMs to produce preference labels for reward learning leads to better performance, in contrast to treating them as reward functions to produce raw reward scores.
@@ V-B Explicit Reward Model
To help robots learn low-level actions, reference [117] harnessed the code generation ability of LLMs to define the lower-level reward parameters based on high-level instructions. Using such a reward design paradigm, their work bridges the gap between high-level language instructions to low-level robot actions and can reliably tackle 90% of the designed tasks compared to the 50% of the baseline. Motivated by the capability of LLM for self-refinement [118], reference [119] proposed a framework with a self-refinement mechanism for automated reward function design, including initial design, evaluation and self-refinement loop. Their results indicate that the LLM-designed reward functions are able to rival or surpass manually designed reward functions. Similarly, Eureka [82] developed a reward optimization algorithm with self-reflection. The algorithm is outlined in Algorithm 2. In each iteration, it uses an environment source code and task description to sample different reward function candidates from a coding LLM. Then the candidates are used to instruct RL training. After training, the results are used to calculate the scores of the reward candidates. Then the best reward function code are selected for reflection, where LLM uses the reasoning capability to progressively improve the reward code. In the experiment, results show that their proposed method can achieve human-level performance on reward design and solve dexterous manipulation tasks that were previously infeasible by manual reward engineering. Another work Text2Reward [120] generated shaped dense reward functions as executable programs based on the environment description. Given the sensitivity of RL training and the ambiguity of language, the RL policy may fail to achieve the goal. Text2Reward addresses the problem by executing the learned policy in the environment, requesting human feedback and refining the reward accordingly.
Algorithm 2 Eureka [82]
0:  Task description $l$, environment code $C$, coding LLM $M_{c}$, evaluation function $E$, initial prompt prompt, optimization iterations $N$, iteration batch size $K$
1:  for $i\leftarrow 1$ to $N$ do
2:     // Sample $K$ reward code candidates from coding LLM $M_{c}$
3:     $R_{1},\dots,R_{K}\leftarrow\texttt{sample}(l,M_{c},C,\texttt{prompt})$
4:     // Evaluate reward candidates
5:     $s_{1}^{i}=E(R_{1}^{i})$, $s_{2}^{i}=E(R_{2}^{i})$, …, $s_{K}^{i}=E(R_{K}^{i})$
6:     // Select the best reward code
7:     ${\mathrm{best}}=\arg\max_{k}(s_{1}^{i},\dots,s_{K}^{i})$
8:     // Reward reflection
9:     $\leftarrow$ prompt:Reflection($R^{i}_{\mathrm{best}},s^{i}_{\mathrm{best}}$)
10:     // Optimize Eureka reward code
11:     if $s^{i}_{\mathrm{best}}>s_{\mathrm{Eureka}}$ then
12:        $R_{\mathrm{Eureka}},s_{\mathrm{Eureka}}\leftarrow(R^{i}_{\mathrm{best}},s^{i}_{\mathrm{best}})$
13:     end if
14:  end for
15:  return  $R_{\mathrm{Eureka}}$
@@ V-C Summarization and Outlook
The use of LLMs as reward designers in RL offers a more natural and efficient way to design complex reward functions. By leveraging natural language processing capabilities, LLMs simplify the traditionally challenging task of reward function design, enhancing both the efficiency and effectiveness of RL algorithms.
However, the inherent biases in LLMs may transfer to the designed reward functions, potentially resulting in suboptimal or harmful behaviors [121]. This presents a potential risk requiring careful consideration. While existing mitigation efforts primarily address biases in demographic, cultural, and political beliefs [122], task-specific biases remain understudied, thus limiting the applicability of LLM-designed reward functions. Toward this end, we list some potential solutions. First, in preference-based RL, when an RL agent optimizes against biased reward models generated by LLMs to predict human preferences, overoptimization and overfitting may occur [123], impeding the learning of the true reward function. Reward function regularization [124] includes the agent preference generated by the value function as a regularization term to mitigate the risk, which helps recover the true underlying reward function. Secondly, human-in-the-loop approaches are another direction to prevent harmful behaviors from designed reward functions. One viable solution is to design an evaluation module where humans can intervene and correct undesired behaviors and reward functions when the agent’s action violates a predefined set of rules. Finally, principles from ensemble learning suggest that combining the reward functions from different LLM models mitigates bias from individual LLMs, leading to improved unbiased performance compared to a single LLM [125].
In the future, we expect advancements in the following areas:
  • Implicit Reward Model: An immediate focus may consider improving how well LLM-generated rewards align with human intentions. This involves refining the quality of language instructions to reduce ambiguities and inaccuracies, ensuring that the reward functions align precisely with human notions of desired behavior. In the longer term, generalization and transferability of LLM-generated rewards across different tasks and environments, especially in complex, high-dimensional visual environments, is also an important direction to explore.
  • Explicit Reward Model: A key limitation in reward code generation is its dependency on pre-trained common-sense knowledge, which can be restrictive for highly specialized tasks not covered in training data. Therefore, short-term goals may include enhancing prompts with detailed, task-specific information and external knowledge. Additionally, the reliance on manually designed templates for motion descriptions limits the adaptability. Looking further ahead, researchers might develop automated or unified processes for designing templates, moving beyond the current limitations of manual motion description templates.
@@ VI LLM as Decision-Maker
LLMs trained on a massive amount of data show impressive results in language understanding tasks [29], instruction-following [126], vision-language navigation [127] and tasks requiring planning and sequential reasoning [83]. Such success motivates researchers to explore the potential of LLMs for decision-making problems. In this section, we divide the role of LLM as 1) the action-maker that generates actions; and 2) the action-guider that instructs the actions.
Figure 5: LLM as a decision-maker. (i) Action-Making: given a $T$-length trajectory $\tau=(\hat{R}_{1},s_{1},a_{1},\dots,\hat{R}_{T},s_{T},a_{T})$ as a sequence of ordered return-to-go $\hat{R}$, action $a$, and states $s$, LLM learns to predict future action $a^{\prime}_{t}$ by minimizing the mean squared error loss $\mathcal{L}=\sum_{t}\left\|a_{t}-a_{t}^{\prime}\right\|_{2}^{2}$. (ii) Action-Guiding: LLM generates a reduced set of action candidates for agents or generates expert actions to regularize RL learning.
VI-A Action-Making
Transformer-based models such as Decision Transformer (DT) [128] have shown great potential in offline RL domain. Instead of using the traditional trial-and-error way, these models treat offline RL as a sequence modeling problem, yielding promising results. As LLM itself is a large-scale Transformer-based model, a natural thought is to leverage the pre-trained power of LLM within this paradigm.
We term this function of LLM as action-making, and identify two typical approaches, as illustrated in Fig. 5 (i). In the first approach (left figure), pre-trained LLM is fine-tuned and then employed for action generation. In the second approach (right figure), goal/instruction along with trajectory are fed to pre-trained LLM. Moreover, a task-specific smaller model is appended after the fine-tuned LLM to facilitate rapid adaptation to diverse tasks.
Pre-trained LLMs outperform basic DT in generalization and sample efficiency, especially for sparse-reward and long-horizon tasks. LLMs’ latent representations, learned from diverse linguistic data, provide valuable prior knowledge for new tasks. This knowledge enables LLMs to solve unseen tasks with less training data, by transferring knowledge from similar tasks and predicting high-reward actions even with sparse feedback. For instance, comparing pre-trained LLM with basic DT, Li et al. [129] reported a 43.6% improvement in out-of-distribution (novel) task completion rates while requiring less training data (e.g., 500 vs 10K). Additionally, Shi et al. [130] demonstrated a 50% performance gain in sparse-reward environments like Kitchen and Reacher2d. For long-horizon tasks, pre-trained representations encode future information, guiding decision-making over extended sequences. For instance, in AntMaze, a long-horizon navigation environment, pre-trained representations yield five times higher scores compared to non-pre-trained counterparts [131].